In preparing dinner you need a cup of very hot water, which you prepare on your electric stove.  You use 40 kJ of electrical energy to heat 300 g of water starting at 20 °C.  What is the final temperature of the water?


1
Expert's answer
2021-03-15T09:26:57-0400

Heat(Q)=40kJ

Mass(m)=300g

Specific heat capacity of water(c)=4.18kJ/Kg


1000g=1kg

300g =?

"\frac {300g×1kg}{1000g}=0.3kg"


But Q=mc"\theta"

40kJ=0.3kg×4.18kJ/kg×(x-20"\degree" C)

40 kJ=1.254x-25.08kJ/kg "\degree" C


"\frac{1.254x}{1.254}=\frac{65.08}{1.254}"


x=51.9"\degree" C